# Deep learning-based sequential data assimilation for chaotic dynamics identifies local instabilities from single state forecasts

### v1

@@ Abstract
We investigate the ability to discover data assimilation (DA) schemes meant for chaotic dynamics with deep learning (DL). The focus is on learning the analysis step of sequential DA, from state trajectories and their observations, using a simple residual convolutional neural network, while assuming the dynamics to be known. Experiments are performed with the Lorenz 96 dynamics, which display spatiotemporal chaos and for which solid benchmarks for DA performance exist. The accuracy of the states obtained from the learned analysis approaches that of the best possibly tuned ensemble Kalman filter (EnKF), and is far better than that of variational DA alternatives. Critically, this can be achieved while propagating even just a single state in the forecast step. We investigate the reason for achieving ensemble filtering accuracy without an ensemble. We diagnose that the analysis scheme actually identifies key dynamical perturbations, mildly aligned with the unstable subspace, from the forecast state alone, without any ensemble-based covariances representation. This reveals that the analysis scheme has learned some multiplicative ergodic theorem associated to the DA process seen as a non-autonomous random dynamical system.
Data assimilation (DA) estimates the state of dynamical systems from sparse and noisy observations, and is used worldwide in numerical weather prediction centers. Accurate DA demands the representation of the time-dependent errors in this state estimate, usually achieved through the propagation of an ensemble of states. Using deep learning, we discover the update step of DA applied to chaotic dynamics. We show that a simple convolutional neural network (CNN) can learn DA, reaching an accuracy as good as that of ensemble-based DA. Crucially, the CNN can achieve this best accuracy with single state forecasts. This is explained by the CNN’s ability to identify local space patterns from this one state, which are used to assess the errors in the analysis. This suggests building a new class of efficient DL-based ensemble-free DA algorithms.
@@ I Introduction
I.1 Sequential data assimilation and accuracy
@@
Sequential data assimilation (DA) is a set of algorithms required to accurately track the trajectory of chaotic dynamics, using a dynamical model and imperfect observations. It is widely used in numerical weather prediction (NWP), and in many areas of climate sciences, (Carrassi et al., 2018) as a suite of both research and operational tools. Classical DA methods are classified into (i) variational methods, such as 3D-Var and 4D-Var, (ii) ensemble-based statistical methods, such as the ensemble Kalman filter (EnKF), and (iii) ensemble variational methods which inherit the assets of the two previous categories.(Asch, Bocquet, and Nodet, 2016) On the one hand, variational methods account for the nonlinearity of models (dynamical model and observation operators), leveraging nonlinear optimization techniques. Ensemble-based methods on the other hand can capture the errors of the day, i.e. time-dependent error statistics, via an ensemble meant to diagnose sample error statistics. Those are key properties that drive the performance of these DA methods in mildly nonlinear chaotic models. For low-order, chaotic dynamics such as the celebrated Lorenz 96 (L96) model,(Lorenz and Emanuel, 1998) the EnKF significantly outperforms 3D-Var, or a moderately-long window 4D-Var in terms of accuracy, owing to its dynamical representation of the errors. This has been emphasized and illustrated in twin experiments. (Bocquet and Sakov, 2013) In fact, current implementations of 4D-Var in NWP centers, incorporate a forecast ensemble so as to capture the errors-of-the-day.(Raynaud, Berre, and Desroziers, 2012; Bonavita, Isaksen, and Hólm, 2012) However, in high-dimensional models, these ensemble-based error statistics must necessarily be regularized using techniques known as localization and possibly inflation.(Evensen, 2009)
I.2 Time-dependent error statistics and dynamical systems
With a focus on the time-dependent error statistics of sequential DA, it has been conjectured (Carrassi et al., 2008a; Palatella, Carrassi, and Trevisan, 2013) then proven (Gurumoorthy et al., 2017; Bocquet et al., 2017; Crisan and Ghil, 2023) that for linear dynamics and when localization is unnecessary, the forecast and analysis error covariance matrices of the EnKF are confined to the unstable-neutral subspace, denoted as ${\mathcal{U}}$ from now on, of the dynamics. This subspace is spanned by the covariant Lyapunov vectors associated to non-negative Lyapunov exponents.(Legras and Vautard, 1996) It is precisely when the ensemble size is smaller than the dimension of this subspace that localization is required to avoid divergence of the EnKF. Deviating from linear dynamics turns those exact results into approximations, for which these findings were nonetheless numerically confirmed.(Bocquet and Carrassi, 2017; Chen, Carrassi, and Lucarini, 2021; Carrassi et al., 2022)
I.3 End-to-end data assimilation and deep learning
@@
It was first proposed to learn DA analysis through DL from the data produced by existing DA schemes. (Härter and de Campos Velho, 2012; Cintra and de Campos Velho, 2018) One can alternatively replace the solver of a 4D-Var over a long DA window by a DL operator that would learn the outcome of the 4D-Var cost function minimization.(Fablet et al., 2021; Frerix et al., 2021; Lafon, Fablet, and Naveau, 2023; Filoche et al., 2023; Keller and Potthast, 2024) However, the latter approaches do not consider cycling sequential DA, the focus of the present paper. A systematic, formal Bayesian view on the use of DL in the critical components of sequential DA has been proposed (Boudier et al., 2023) and called data assimilation network (DAN). In the present paper, a simplified variant of this DAN concept is used. As far as ensemble and Kalman-related DA methods are concerned, it has been proposed to learn their Kalman gain, (Hoang, De Mey, and Talagrand, 1994; Hoang et al., 1998) or parameters thereof, possibly relying on an auto-differentiable implementation of the (En)KF. (Haarnoja et al., 2016; Chen, Sanz-Alonso, and Willett, 2022; Luk et al., 2024) A step further, it was also proposed to learn the fully analysis operator using (self-)supervision.(McCabe and Brown, 2021; Boudier et al., 2023) Finally, bypassing the need for dynamical models and DA schemes altogether, DL-based end-to-end methods aim at estimating states of the system from the observations only,(McNally et al., 2024; Vaughan et al., 2024) but with a focus on the feasibility of such endeavor for now.
I.4 Objective
In this paper, the forecast model is assumed to be known so as to disentangle intricate interactions of learning simultaneously the DA operators and the dynamics altogether. Our objective is to learn the analysis step of a sequential DA scheme meant for chaotic dynamics from a long trajectory of the dynamical system and the associated set of noisy, potentially partial observations. Supervised learning is used although self-supervised learning(Bocquet et al., 2019; Brajard et al., 2020; McCabe and Brown, 2021) could have been used instead but is more challenging and rather disconnected from the main results of this paper. The (incremental) analysis CNN operator will be referred to as $a_{\bm{\theta}}$, while the full resulting DA scheme will be called DAN.
We will show that the learned $a_{\bm{\theta}}$ directly discovers and utilizes a fine knowledge of the dynamics, as opposed to agnostic classical DA. The nature of these fine dynamical structures learned through $a_{\bm{\theta}}$ will then be discussed and interpreted, together with key implications for the development of modern DA methods.
@@ II Experimental setup
We use a twin experiment setup where the states of a trajectory of the dynamics, ${\mathbf{x}}_{k}^{\mathrm{t}}\in\mathbb{R}^{N_{\mathrm{x}}}$ at equally spaced times $t_{k}$ for $k=1\ldots N_{\mathrm{c}}$, are generated from the known dynamics; the superscript “t” stands for truth. A set of observation vectors $\mathbf{y}_{k}\in\mathbb{R}^{N_{\mathrm{y},k}}$ are obtained from these states ${\mathbf{x}}_{k}^{\mathrm{t}}$ via an observation operator $\mathcal{H}_{k}$, and perturbed by a white-in-time Gaussian noise ${\bm{\varepsilon}}_{k}$ of mean $\mathbf{0}$ and covariance matrix $\mathbf{R}_{k}\in\mathbb{R}^{N_{\mathrm{y},k}\times N_{\mathrm{y},k}}$:
$\mathbf{y}_{k}=\mathcal{H}_{k}({\mathbf{x}}_{k}^{\mathrm{t}})+{\bm{\varepsilon}}_{k},\qquad{\bm{\varepsilon}}_{k}\sim N(\mathbf{0},\mathbf{R}_{k}).$ (1)
@@ II.1 Analysis operator and its neural network representation
Let us define a sequential DA scheme, based on an analysis and forecast (also called background) ensemble. The $i$-th members of the analysis and forecast ensembles at time $t_{k}$ are noted ${\mathbf{x}}_{k}^{\mathrm{a},i}$ and ${\mathbf{x}}_{k}^{\mathrm{f},i}$, respectively. Denoting ${\mathcal{S}}_{\mathrm{e}}=1\ldots N_{\mathrm{e}}$, the corresponding analysis and forecast ensembles are ${\mathbf{E}}_{k}^{\mathrm{a}}=\left\{{\mathbf{x}}_{k}^{\mathrm{a},i}\right\}_{i\in{\mathcal{S}}_{\mathrm{e}}}\subset{\mathcal{E}}^{\mathrm{e}}$, ${\mathbf{E}}_{k}^{\mathrm{f}}=\left\{{\mathbf{x}}_{k}^{\mathrm{f},i}\right\}_{i\in{\mathcal{S}}_{\mathrm{e}}}\subset{\mathcal{E}}^{\mathrm{e}}$, respectively, where ${\mathcal{E}}^{\mathrm{e}}=\mathbb{R}^{N_{\mathrm{e}}\times N_{\mathrm{x}}}$. The initial ensemble ${\mathbf{E}}_{0}^{\mathrm{f}}$ is obtained from perturbing a random state on the attractor of the dynamics.
The analysis step of the DA scheme is given by the (incremental) analysis operator $a_{\bm{\theta}}$, which depends on a set of CNN weights and biases, a vector $\bm{\theta}$,
@@
where ${\bm{\delta}}_{k}$, the innovation at time $t_{k}$, is defined by
@@
$\mathbf{H}_{k}$ is the tangent linear operator of $\mathcal{H}_{k}$ but any arbitrary injective operator from $\mathbb{R}^{N_{\mathrm{y},k}}$ to $\mathbb{R}^{N_{\mathrm{x}}}$ could be chosen instead. The DA forecast step propagates the analysis ensemble,
${\mathbf{E}}_{k+1}^{\mathrm{f}}={\mathcal{M}}\left({\mathbf{E}}_{k}^{\mathrm{a}}\right),$ (3)
member-wise; $\mathcal{M}$ is the resolvent over $t_{t+1}-t_{k}$ of the autonomous, i.e. time-independent, dynamics.
In this paper, we choose $a_{\bm{\theta}}$ to be a simple residual CNN, whose architecture is detailed in Appendix A. The internal convolutional layers are chosen to have $N_{\mathrm{f}}$ channels. Hence, the internal state of the CNN consists of $N_{\mathrm{f}}$ copies of the latent space which we simply chose to be isomorphic to the state space $\mathbb{R}^{N_{\mathrm{x}}}$. Furthermore, the encoder and decoder from state space to latent space and back, are chosen to coincide with the identity.
@@ II.2 Training scheme
Towards efficiently learning an optimal $a_{\bm{\theta}}$, we consider $N_{\mathrm{r}}$ such DA runs, based on as many independent concurrent trajectories of the dynamics and as many sequences of observation vectors. The DA runs are hence specified by ${\mathbf{E}}_{k}^{\mathrm{a},r}=\left\{{\mathbf{x}}_{k}^{\mathrm{a},i,r}\right\}_{i\in{\mathcal{S}}_{\mathrm{e}}}$ and ${\mathbf{E}}_{k}^{\mathrm{f},r}$ for $r=1\ldots N_{\mathrm{r}}$ and, being iterates through $N_{\mathrm{c}}$ cycles, they depend on $\bm{\theta}$, except for the set of initial conditions ${\mathbf{E}}_{0}^{\mathrm{f},r}$. In order to learn an optimal $a_{\bm{\theta}}$, a loss function is defined:
@@
It must be pointed out that a proper sequential DA process when seen as a dynamical system, is stable. (Carrassi et al., 2008b, c) Hence, after a rough starting phase in the training, the learned $a_{\bm{\theta}}$ should yield a numerically stable prediction-assimilation dynamical system. In particular, this is likely to avoid exploding gradients. By contrast, the task of learning a DL emulator of the dynamics over many consecutive time steps fails because of the unstable nature of the chaotic dynamics.
The $N_{\mathrm{r}}$ trajectories are dispatched into a training and validation dataset with a $90\%-10\%$ ratio. Moreover, the testing dataset is generated from an independently generated trajectory, long enough to yield converged statistics. In the test stage, the DAN scheme is used within a twin experiment using the trajectories and resulting observations of the testing dataset. Its performance is assessed using the time-averaged root mean square error (aRMSE) against the truth.

### v2

# Accurate deep learning-based filtering for chaotic dynamics by identifying instabilities without an ensemble

@@ Abstract
We investigate the ability to discover data assimilation (DA) schemes meant for chaotic dynamics with deep learning. The focus is on learning the analysis step of sequential DA, from state trajectories and their observations, using a simple residual convolutional neural network, while assuming the dynamics to be known. Experiments are performed with the Lorenz 96 dynamics, which display spatiotemporal chaos and for which solid benchmarks for DA performance exist. The accuracy of the states obtained from the learned analysis approaches that of the best possibly tuned ensemble Kalman filter, and is far better than that of variational DA alternatives. Critically, this can be achieved while propagating even just a single state in the forecast step. We investigate the reason for achieving ensemble filtering accuracy without an ensemble. We diagnose that the analysis scheme actually identifies key dynamical perturbations, mildly aligned with the unstable subspace, from the forecast state alone, without any ensemble-based covariances representation. This reveals that the analysis scheme has learned some multiplicative ergodic theorem associated to the DA process seen as a non-autonomous random dynamical system.
Data assimilation (DA) estimates the state of dynamical systems from sparse and noisy observations, and is used worldwide in numerical weather prediction centers. Accurate DA demands the representation of the time-dependent errors in this state estimate, usually achieved through the propagation of an ensemble of states. Using deep learning, we discover the update step of DA applied to chaotic dynamics. We show that a simple convolutional neural network (CNN) can learn DA, reaching an accuracy as good as that of ensemble-based DA. Crucially, the CNN can achieve this best accuracy with single state forecasts. This is explained by the CNN’s ability to identify local space patterns from this one state, which are used to assess the errors in the analysis. This suggests building a new class of efficient deep learning-based ensemble-free DA algorithms.
@@ I Introduction
I.1 Context and problem
In a simplified but quintessential framework, the goal of data assimilation (DA), and in particular filtering algorithms, is to accurately estimate states ${\mathbf{x}}_{k}^{\mathrm{t}}\in\mathbb{R}^{N_{\mathrm{x}}}$, where “t” stands for truth, at equally spaced times $\tau_{k}$ for $k=0,\ldots,K$ along a trajectory of a dynamical system. Hence, they are related by
${\mathbf{x}}_{k+1}^{\mathrm{t}}={\mathcal{M}}\left({\mathbf{x}}_{k}^{\mathrm{t}}\right),$ (1a)
where $\mathcal{M}$ is the resolvent over $\tau_{k+1}-\tau_{k}$ of known autonomous, i.e. time-independent, dynamics. Such goal is achieved from the knowledge of the dynamics $\mathcal{M}$ and of observation vectors $\mathbf{y}_{k}\in\mathbb{R}^{N_{\mathrm{y},k}}$ obtained from the non-accessible states ${\mathbf{x}}_{k}^{\mathrm{t}}$ via observation operators $\mathcal{H}_{k}$, and perturbed by a white-in-time Gaussian noise ${\bm{\varepsilon}}_{k}$ of mean $\mathbf{0}$ and covariance matrix $\mathbf{R}_{k}$:
$\mathbf{y}_{k}=\mathcal{H}_{k}({\mathbf{x}}_{k}^{\mathrm{t}})+{\bm{\varepsilon}}_{k},\qquad{\bm{\varepsilon}}_{k}\sim N(\mathbf{0},\mathbf{R}_{k}).$ (1b)
Applied to chaotic hence dynamically unstable dynamics, sequential (in time) algorithms must be used.(Asch, Bocquet, and Nodet, 2016; Carrassi et al., 2018) They alternate an analysis step which, from the newly acquired observation vector $\mathbf{y}_{k}$ in Eq. (1b) and the current estimate of the state ${\mathbf{x}}_{k}^{\mathrm{f}}$, provides an updated optimal estimate of the state ${\mathbf{x}}_{k}^{\mathrm{a}}$ called the analysis. The subsequent state estimate ${\mathbf{x}}_{k+1}^{\mathrm{f}}$ stems from the forecast step which relies on Eq. (1a). The estimates in both steps can either be deterministic or probabilistic, often leveraging an ensemble in the latter case. Such sequential DA is widely used in numerical weather prediction (NWP), and in many areas of climate sciences,(Carrassi et al., 2018) as a suite of both research and operational tools.
@@ I.1 Context and problem
Classical DA methods are classified into (i) variational methods, such as 3D-Var and 4D-Var, (ii) ensemble-based statistical methods, such as the ensemble Kalman filter (EnKF), and (iii) ensemble variational methods which inherit the assets of the two previous categories.(Asch, Bocquet, and Nodet, 2016) On the one hand, variational methods account for the nonlinearity of models (dynamical model and observation operators), leveraging nonlinear optimization techniques. Ensemble-based methods on the other hand can capture the errors of the day, i.e. time-dependent error statistics, via an ensemble meant to diagnose sample error statistics. Those are key properties that drive the performance of these DA methods in mildly nonlinear chaotic models. For low-order, chaotic dynamics such as the celebrated Lorenz 96 (L96) model,(Lorenz and Emanuel, 1998) the EnKF significantly outperforms 3D-Var, or a moderately-long window 4D-Var in terms of accuracy, owing to its dynamical representation of the errors. This has been emphasized and illustrated in twin experiments.(Bocquet and Sakov, 2013) In fact, current implementations of 4D-Var in NWP centers incorporate a forecast ensemble so as to capture the errors-of-the-day.(Raynaud, Berre, and Desroziers, 2012; Bonavita, Isaksen, and Hólm, 2012) However, in high-dimensional models, these ensemble-based error statistics must necessarily be regularized using techniques known as localization and possibly inflation.(Evensen, 2009) With a focus on the time-dependent error statistics of sequential DA, it has been conjectured (Carrassi et al., 2008a; Palatella, Carrassi, and Trevisan, 2013) then proven (Gurumoorthy et al., 2017; Bocquet et al., 2017; Crisan and Ghil, 2023) that for linear dynamics and when localization is unnecessary, the forecast and analysis error covariance matrices of the EnKF are confined to the unstable-neutral subspace, denoted as ${\mathcal{U}}$ from now on, of the dynamics. This subspace is spanned by the covariant Lyapunov vectors associated to non-negative Lyapunov exponents.(Legras and Vautard, 1996) It is precisely when the ensemble size is smaller than the dimension of this subspace that localization is required to avoid divergence of the EnKF. Deviating from linear dynamics turns those exact results into approximations, for which these findings were nonetheless numerically confirmed.(Bocquet and Carrassi, 2017; Chen, Carrassi, and Lucarini, 2021; Carrassi et al., 2022)
@@
It was first proposed to learn DA analysis through DL from the data produced by existing DA schemes. (Härter and de Campos Velho, 2012; Cintra and de Campos Velho, 2018) One can alternatively replace the solver of a 4D-Var over a long DA window by a DL operator that would learn the outcome of the 4D-Var cost function minimization.(Fablet et al., 2021; Frerix et al., 2021; Lafon, Fablet, and Naveau, 2023; Filoche et al., 2023; Keller and Potthast, 2024) However, the latter approaches do not consider cycling sequential DA, the focus of the present paper. A systematic, formal Bayesian view on the use of DL in the critical components of sequential DA has been proposed (Boudier et al., 2023) and called data assimilation network (DAN). In the present paper, a simplified variant of this DAN concept is used. As far as ensemble and Kalman-related DA methods are concerned, it has been proposed to learn their Kalman gain, (Hoang, De Mey, and Talagrand, 1994; Hoang et al., 1998) or parameters thereof, possibly relying on an auto-differentiable implementation of the (En)KF. (Haarnoja et al., 2016; Chen, Sanz-Alonso, and Willett, 2022; Luk et al., 2024) As a step further, it was also proposed to learn the full analysis operator using (self-)supervision.(McCabe and Brown, 2021; Boudier et al., 2023) Finally, bypassing the need for dynamical models and DA schemes altogether, DL-based end-to-end methods aim at estimating states of the system from the observations only,(McNally et al., 2024; Vaughan et al., 2024) yet so far with a focus on the feasibility of such endeavor.
I.2 Objectives
In this paper, the forecast model in Eq. (1a) is assumed to be known so as to avoid intricate interactions when learning the DA operators and the dynamics simultaneously.
Our objective is to learn the analysis operator of a sequential DA scheme meant for chaotic dynamics from a long trajectory of the dynamical system and the associated set of noisy, possibly sparse observations. Hence, it stands out from past studies that exploited DL to learn the dynamics, possibly using DA. (Bocquet et al., 2019; Brajard et al., 2020; Bocquet et al., 2020; Brajard et al., 2021; Liu et al., 2022; Wang et al., 2024) The resulting DL-based analysis operator will be referred to as $a_{\bm{\theta}}$, while the full resulting DA scheme will be called DAN.
We will first explain how to learn such analysis operator from DL. It will then be shown numerically that $a_{\bm{\theta}}$ can surprisingly perform as accurately as a well optimized EnKF, even without using an ensemble which strongly contrasts with the common beliefs in methodological DA. To interpret this result, we will show using innovative concepts based on a Taylor expansion of the learned $a_{\bm{\theta}}$, that $a_{\bm{\theta}}$ directly discovers and utilizes a fine knowledge of the dynamics, as opposed to agnostic classical DA. The nature of these dynamical structures learned through $a_{\bm{\theta}}$ will then be discussed and interpreted.
@@ II Experimental setup
With the goal to learn an analysis operator $a_{\bm{\theta}}$ as a key step of a filtering DA scheme for chaotic dynamics, we build a twin experiment within the framework offered by Eqs. (1).
@@ II.1 Analysis operator and its neural network representation
Let us define a filtering DA scheme, based on an analysis and forecast ensemble. The $i$-th members of the analysis and forecast ensembles at time $\tau_{k}$ are noted ${\mathbf{x}}_{k}^{\mathrm{a},i}$ and ${\mathbf{x}}_{k}^{\mathrm{f},i}$, respectively. Denoting ${\mathcal{S}}_{\mathrm{e}}=1,\ldots,N_{\mathrm{e}}$, the corresponding analysis and forecast ensembles are ${\mathbf{E}}_{k}^{\mathrm{a}}=\left\{{\mathbf{x}}_{k}^{\mathrm{a},i}\right\}_{i\in{\mathcal{S}}_{\mathrm{e}}}\subset{\mathcal{E}}^{\mathrm{e}}$, ${\mathbf{E}}_{k}^{\mathrm{f}}=\left\{{\mathbf{x}}_{k}^{\mathrm{f},i}\right\}_{i\in{\mathcal{S}}_{\mathrm{e}}}\subset{\mathcal{E}}^{\mathrm{e}}$, respectively, where ${\mathcal{E}}^{\mathrm{e}}=\mathbb{R}^{N_{\mathrm{e}}\times N_{\mathrm{x}}}$. The initial ensemble ${\mathbf{E}}_{0}^{\mathrm{f}}$ is obtained from perturbing a random state on the attractor of the dynamics.
The analysis step of the DA scheme is given by the (incremental) analysis operator $a_{\bm{\theta}}$, which depends on a set of neural network weights and biases, a vector $\bm{\theta}$,
@@
where ${\bm{\delta}}_{k}$, the innovation at time $\tau_{k}$, is defined by
@@
$\mathbf{H}_{k}$ is the tangent linear operator of $\mathcal{H}_{k}$ but any arbitrary injective operator from $\mathbb{R}^{N_{\mathrm{y},k}}$ to $\mathbb{R}^{N_{\mathrm{x}}}$ could be chosen instead. The DA forecast step propagates the analysis ensemble, member-wise:
${\mathbf{E}}_{k+1}^{\mathrm{f}}={\mathcal{M}}\left({\mathbf{E}}_{k}^{\mathrm{a}}\right).$ (3)
We choose $a_{\bm{\theta}}$ to have a simple residual convolutional neural network (CNN) architecture. A schematic of the CNN architecture is displayed in Fig. 1.
Figure 1: Architecture of the residual convolutional network, where $N_{\mathrm{b}}=2$, $N_{\mathrm{sb}}=3$. $\mathrm{conv}_{N_{1},N_{2},f}$ is a generic one-dimensional convolutional layer of dimension $N_{1}$, with $N_{2}$ filters of kernel size $f$. See text for more details.
It begins with an initial convolution that takes $N_{\mathrm{e}}+1$ channels as inputs and, with $N_{\mathrm{f}}$ filters, outputs $N_{\mathrm{f}}$ channels. This initial layer is followed by $N_{\mathrm{b}}$ residual blocks. Each one of these blocks is a succession of $N_{\mathrm{sb}}$ sub-blocks. Each subblock is made of: (i) a convolutional layer with $N_{\mathrm{f}}$ channels as inputs, which has $N_{\mathrm{f}}$ filters and a kernel size $f_{\mathrm{v}}$ for each of its filter, (ii) a batch normalization layer, and (iii) an activation function chosen to be mish. (Misra, 2019) The CNN ends with a final convolutional layer that takes $N_{\mathrm{f}}$ channels as inputs and, with $N_{\mathrm{e}}$ filters, outputs $N_{\mathrm{e}}$ channels. The kernel size of the initial and final channels is $f_{\mathrm{v}}$. Hence, the internal state of the CNN consists of $N_{\mathrm{f}}$ copies of the latent space which we simply choose to be isomorphic to the state space $\mathbb{R}^{N_{\mathrm{x}}}$. Furthermore, the encoder and decoder from state space to latent space and back are chosen to coincide with the identity. We have also tested a depth-separable architecture for this residual CNN, with a number of parameters roughly divided by $3$, yielding an accuracy almost as good but longer training times. Note that the fundamental results reported in this paper are agnostic to the details of the architecture: this CNN is a mere functional tool to learn an optimal $a_{\bm{\theta}}$.
@@ II.2 Training scheme
Towards efficiently learning an optimal $a_{\bm{\theta}}$, we consider $N_{\mathrm{r}}$ such DA runs, based on as many independent concurrent trajectories of the dynamics and as many sequences of observation vectors. Hence, the DA runs are specified by ${\mathbf{E}}_{k}^{\mathrm{a},r}=\left\{{\mathbf{x}}_{k}^{\mathrm{a},i,r}\right\}_{i\in{\mathcal{S}}_{\mathrm{e}}}$ and ${\mathbf{E}}_{k}^{\mathrm{f},r}$ for $r=1,\ldots,N_{\mathrm{r}}$ and, being iterates through $N_{\mathrm{c}}$ cycles, they depend on $\bm{\theta}$ except for the set of initial conditions ${\mathbf{E}}_{0}^{\mathrm{f},r}$. In order to learn an optimal $a_{\bm{\theta}}$, a loss function is defined:
@@
It must be pointed out that a successful sequential DA process, when seen as a dynamical system, is stable. (Carrassi et al., 2008b, c) Hence, after a rough starting phase in the training, the learned $a_{\bm{\theta}}$ should yield a numerically stable prediction-assimilation dynamical system. In particular, this is likely to avoid exploding gradients. By contrast, the task of learning a DL emulator of the dynamics over many consecutive time steps often fails because of the unstable nature of the chaotic dynamics.
The $N_{\mathrm{r}}$ trajectories are dispatched into a training and validation dataset with a $90\%-10\%$ ratio. Overfitting is prevented by an early stopping of the minimization based on the validation score, tantamount to regularization.(Goodfellow, Bengio, and Courville, 2016) Moreover, the testing dataset stems from an independently generated trajectory, long enough to yield converged statistics. In the test stage, the DAN scheme is used within a twin experiment using the trajectories and resulting observations of the testing dataset. Its performance is assessed from a single scalar score using the time-averaged root mean square error (aRMSE) of the analysis against the truth:
$\mathrm{aRMSE}=\frac{1}{K\sqrt{N_{\mathrm{x}}}}\sum_{k=1}^{K}\left\|{\mathbf{x}}_{k}^{\mathrm{t}}-\bar{\mathbf{x}}_{k}^{\mathrm{a}}\right\|,$ (5)
which, in a cycled DA context, is a reliable indicator of the overall performance of the scheme, whatever its purpose.